OpenStudy (anonymous):

Medals shall be given to the best response :) Please help, I have difficulty solving such questions. A rectangular garden, 21m^2 in area, will be fenced to keep out rabbits and skunks. Find the dimensions that will require the least amount of fencing if a barn already protects one side of the garden.

OpenStudy (mww):

First construct an equation(s) that model the problem. You can draw a rectangle.  Let one of dimensions be x and one y (doesn't matter which one) The perimeter is P = x + y + x (given one side, y, is the barn that already encloses the garden) \[P = 2x+y\] We also know the area is 21 sqm. The area of the rectangle is A = xy \[A = xy = 21, y=21/x\] Now we can substitue y = 21/x into the perimeter formula. \[P = 2x + y = 2x+\frac{ 21 }{ x }\] All we need to do is find the first derivative and solve for zero as this tells us possible maxima or minima. \[\frac{ dP }{ dx } = 2 - \frac{ 21 }{ x^2 } = \frac{ 2x^2 - 21 }{ x^2 } = 0\] So we solve \[2x^2 - 21 = 0 -> 2x^2 = 21 -> x^2 = 21/2 -> x = \sqrt{\frac{ 21 }{ 2 }}\] taking x to be positive as it is a length. Now you can confirm second differentiation that it is a minimum or check with first derivative around the point on both sides of the turning point. Then substitute into any formula involving x to find y.

OpenStudy (anonymous):

Thanks, but we are using methods in the advanced functions course. We can't use derivatives.

OpenStudy (mww):

Are you allowed to sketch? If you do, you should see that adding the graphs of y=2x and y =21/x together (which add to give P, the perimeter), the minimum occurs when they intersect for positive x. So solve for 2x = 21/x and you will find such x.

OpenStudy (anonymous):

Okay let me try it out

OpenStudy (anonymous):

Could you explain the last step to me again. I don't understand it.

OpenStudy (anonymous):

I end up with 2y^2 + 21/y = 0 When I solve that, we will have to take the square root of a negative. Which is impossible.

OpenStudy (mww):

No no this is what I mean. Sketch y =2x and y =21/x (very roughly) on the same axes and y = 2x +21/x The intersection of y=2x and y=21/x is where y =2x +21x will be a minimum. Compare: http://www.wolframalpha.com/input/?i=2x+%3D+21%2Fx http://www.wolframalpha.com/input/?i=2x+%2B+21%2Fx So you solve NOT P = 0, but 2x = 21/x

OpenStudy (anonymous):

Okay. The perimeter = 2x + y Area = y = 21/x I know where the y=21/x graph comes from, where do you get y = 2x?

OpenStudy (mww):

From P =2x +21/x. to draw it you can add ordinates from y = 21x to y = 21/x. It happens so when you add 2x to 21/x where they intersect, you get a minimum.

OpenStudy (mww):

*y=2x to y =21/x

OpenStudy (anonymous):

All i need to know is how you determined that y = 2x, that is the only part bothering me.

OpenStudy (mww):

oh well perimeter is 2x + y right? you also found y = 21/x...substitute into perimeter: P = 2x + y = 2x + 21/x So there's the 2x. All I did was draw 2x and 21/x on the same axes, then added points together to give 2x + 21/x.

OpenStudy (anonymous):

Oh, so you treated 2x and 21/x as two seperate graphs?

OpenStudy (mww):

yes. You don't need to but when you draw them separately but on the same set of axes, and then add points together, you find easily that the minimum occurs when y=2x happens to intersect with y =21/x.

OpenStudy (anonymous):

Okay. Thank you very much :)

OpenStudy (mww):

This might make more sense with all three graphs http://www.wolframalpha.com/input/?i=graph%282x%2C+21%2Fx%2C+2x+%2B+21%2Fx%2C+x%3D-10..10%2C+y%3D+-20..20%29

OpenStudy (anonymous):

Thanks.